OpenStudy (joshoyen):

PLEASE HELP ASAP. BEEN STUCK ON THIS FOR 1 HOUR Solve each equation. Check each solution. \[x+\frac{ 10 }{ x-2 }=\frac{ x^2+3x }{ x-2 }\]

OpenStudy (faiqraees):

multiply each term with x -2

OpenStudy (joshoyen):

oh wait, wth i put the wrong problem. its \[\frac{ 1 }{ 2x+2 }+\frac{ 5 }{ x^2-1 } = \frac{ 1 }{ x-1 }\] sorry bout that

OpenStudy (faiqraees):

okay factor the denominator of second fraction

OpenStudy (joshoyen):

\[\frac{ 5 }{ (x+1)(x-1) }\]

OpenStudy (joshoyen):

not sure

OpenStudy (faiqraees):

Perfect now multiply each term wit (x+1)(x-1)

OpenStudy (joshoyen):

so \[(x+1)(x-1)\frac{ 1 }{ 2x+2 } ?\]

OpenStudy (faiqraees):

\[ (x+1)(x-1)\frac{1}{2(x+1)} +(x+1)(x-1)\frac{ 5 }{ (x+1)(x-1) }= (x+1)(x-1)\frac{ 1 }{ x-1 } \]

OpenStudy (faiqraees):

now simplify them

OpenStudy (joshoyen):

Hmm.. not sure on how to do this part

OpenStudy (faiqraees):

Cancel out the like terms in numerator and denominator

OpenStudy (joshoyen):

Is it \[(x-1)\frac{ 1 }{ 2 }+\frac{ 5 }{ 1 }+(x+1)\frac{ 1 }{ 1 }\] ?

OpenStudy (anonymous):

x = 7 \[\frac{5}{x^2-1}+\frac{1}{2 x+2}-\frac{1}{x-1}=0 \]\[\frac{7-x}{2 (x-1) (x+1)}=0 \]x = 7

OpenStudy (joshoyen):

ohhh okay, thanks guys!